# An engineering guide to superconducting quantum circuit shielding

### v1

@@ III.5 Drive line filtering (XY control)
For the SQCs driving, special drive lines are utilized. In some implementations, the SQC driving is carried out through a readout line. However, in this review, we will consider an individual drive lines separated from readout, and analyze their filtration. At the SQCs drive line, the following filtering configurations are commonly used: LPF, IR, or LPF+IR (see Fig. 19). More complex combination (e.g. LPF+IRx2) is rarely used. BPF is also less frequently utilized on this line. These filters are often placed at the lower stage of the fridge, but they can also be placed at the 100 $mK$ stage or at room temperature.
III.6 Flux line filtering (Z control)
@@
The flux line filtering is usually based on application of LPF and IR filters. For this line, as for the drive line, we consider a separated flux line (in some implementations the magnetic flux can be manipulated by a special coil surrounding the sample holder). The most commonly used configurations of the filters on the flux line are LPF+IR, LPF or IR+LPF. Generally, the flux line filters are placed at the lower stage of the fridge, at the same time an LPF can be placed on other stages of the cryostat.
@@ V SUMMARY AND OUTLOOK
In this review, we have described the requirements for the shielding system to protect the SQC from external IR radiation and stray electromagnetic fields. We have considered a variety of existing shielding systems and compare them to our requirements. We have also considered different techniques for shielding system effectiveness evaluation.
@@ V.1 Recommendations for shielding and filtering
    The shielding system should consist of at least three shields: an IR shield, a superconducting shield, and a ${\mu}$-metal one.
@@
    All the shields should be light-tight. It is necessary to prevent the penetration of IR radiation inside the shield. For superconducting and ${\mu}$-metal shields, it allows for the magnetic field equalising in the entire shield volume.
@@
    For creating IR shielding with a high efficiency, it should consist of a base with a high thermal conductivity (copper or superconductor coated copper). The shield should have a light-tight lid sealed by the indium wire. The inner surface of IR shield should be coated by the mixture of absorbing coating composed of Stycast epoxy and SiC particles of about 1 $mm$. The shield should be placed as close to the SQC as possible.
@@
    For protection from magnetic fields, it is recommended to use a combination of two shields, i.e. an internal superconducting shield and an external ${\mu}$-metal shield. Such shields arrangement allows for the lowest possible level of stray magnetic field inside the shields.
@@
  8. 8.
    The placement of magnetic components of the cryogenic setup (e.g., isolators, circulators) should be done outside the shielding. A special design features, such as waveguides on the lid, can additionally reduce the influence of the magnetic field on the quantum circuit. Besides, placing the dilution refrigerator inside a shielded laboratory or in a deep underground location can reduce the influence of the cosmic rays.
@@
    The sample holder serves as the first layer of shielding. The recommendations for its fabrication are similar to those for IR shielding (see point 4 of current recommendations list). The design of the sample holder should be selected based on the SQC design and the wiring method to the cryogenic setup.
@@
    The best protection of the superconducting quantum circuit from stray electromagnetic fields and IR radiation can be achieved by combining the shielding and microwave filters. To effectively filter the input line, both RF and MW filters and IR ones should be used, specifically BPF+IR or HPF+LPF+IR. Eccosorb CR-110 epoxy should be chosen as a filler for the IR filters. The IR filters should be placed inside the shield thermalized on the lower stage of dilution refrigerator, while the RF and MW filters don’t have such requirement. The output readout line should be filtered in the same manner as the input one. For drive and flux lines, LPF+IR should be used, because it protects from IR radiation propagating along the microwave lines.

### v2

@@ III.3 Drive line filtering (XY control)
For the SQCs driving, special drive lines are utilized. In some implementations, the SQC driving is carried out through a readout line. However, in this review, we will consider an individual drive lines separated from readout.
LPFs or BPFs are essential on qubit drive lines to suppress parasitic harmonics generated during frequency up-conversion. These filters can be placed either at room temperature or at the base of the cryostat. For room-temperature implementation, an isolator must be positioned upstream of the filter to prevent standing wave formation and subsequent control pulse distortionLi et al. (2023). In contrast, placement at the lowest cryostat stage utilizes attenuators to inherently suppress standing wavesKelly (2015), rendering additional isolation unnecessary.
Additional IR filtering is required for the qubit drive line, as standard microwave LPFs and BPFs provide no specified attenuation outside their operational bandwidth. For instance, the Marki FLP-0750 filter exhibits effective performance only up to 20 $GHz$Marki Microwave . The IR filter must be positioned between attenuators to ensure proper impedance matchingKrinner et al. (2019). Furthermore, it requires enclosure within a shield anchored to the cryostat’s lower stage to prevent infrared radiation leakage from warmer stages into the post-filter segment of the drive lineSerniak (2019).
III.4 Flux line filtering (Z control)
@@
The flux line filtering is usually based on application of LPF and IR filters. For this line, as for the drive line, we consider a separated flux line (in some implementations the magnetic flux can be manipulated by a special coil surrounding the sample holder).
LPFs are essential on qubit flux lines to suppress noise bandwidth contributions from both warmer cryogenic stages and control electronics. The filter’s cutoff frequency must be selected according to the required flux pulse duration - for instance, a minimum 200 $MHz$ cutoff is necessary to support fast two-qubit operations with 15 $ns$ pulsesFoxen et al. (2020). In practice, 500 $MHz$ filters are commonly usedArute et al. (2019) as a trade-off between fast two-qubit operations and the noise bandwidth.
IR filters are essential on flux lines to protect against Cooper-pair-breaking radiation. Since LPFs operate via signal reflection (rejecting frequencies above cutoff), this reflection mechanism creates standing waves between the LPF and qubit at operational frequencies, inducing qubit relaxationKlimov et al. (2018). To suppress these parasitic modes, an IR filter must be inserted between the LPF and qubitIvanov et al. (2023). This IR filter must exhibit minimal reflection at qubit frequencies (less than -15 $dB$). For these applications, Eccosorb CR-124 serves as an optimal IR absorber due to its frequency-selective properties: low absorption (< 0.5 $dB/cm$) below 500 $MHz$ and high absorption ($\sim$ 40 $dB/cm$) at typical qubit frequenciesEMERSON and CUMING MICROWAVE PRODUCTS . Furthermore, Eccosorb CR-124-based filters enable combined flux and drive line implementations Li et al. (2025); Manenti et al. (2021), as this material simultaneously provides sufficient IR attenuation for flux line protection, and adequate microwave attenuation for drive line.
III.5 Filtering of readout line input
Typically, in cryogenic setups one Córcoles et al. (2015); Wu et al. (2021); Buks et al. (2020); Martin et al. (2020); Lachance-Quirion et al. (2017) or two filters Gyenis et al. (2021); Schneider et al. (2019); Opremcak et al. (2021) are used at the input readout line. Less frequently, three filters are installed in a row Chen (2018); Colless et al. (2018); Grünhaupt et al. (2019). Microwave and IR filters can be used both individually Schwartz (2016); Yan et al. (2016); Grünhaupt et al. (2017); Andersen et al. (2019) and together Zhang (2018); Lisenfeld et al. (2019); Leonard et al. (2019).
In most measurement circuits, filters are located outside the shielding Jin et al. (2015); Arute et al. (2019); Yeh et al. (2017); Jr. (2018); Kudra et al. (2020); Wilen et al. (2021). Less frequently, they may be located both inside and outside the shielding Burnett et al. (2019); Premkumar et al. (2021); Gao (2018); Mundhada (2020). In very rare cases, the filters are only located inside the shielding Kreikebaum et al. (2016), mounted on the lid of the shield Hazard (2019); Barends et al. (2011), or just prior to the sample holder Cooper (2013). In Serniak (2019) it is indicated that the location of an IR filter inside the shielding can reduce the quasiparticles generation in a qubit. Thus, it is recommended to place the IR filter within a metal shield thermalized on the base stage of the dilution refrigerator. For microwave filters there is no special requirements for positioning in the measurement circuit. Thus, they are typically located outside the shielding. The filtering requirements for readout line input are generally equivalent to those for qubit drive lines, necessitating comparable filter specifications and implementation strategies.
III.6 Filtering of readout line output
The readout line output requires minimal signal attenuation to preserve the signal-to-noise ratio (SNR), necessitating the use of non-dissipative filters (low-pass or band-pass filters) and isolators. This configuration provides three critical benefits: firstly, suppression of thermal noise propagation from warmer cryostat stages; secondly, prevention of signal reflection from the output line back to the qubit; and thirdly, isolation of parametric amplifier pump leakage. For IR radiation mitigation, either compact IR filters (to minimize output signal attenuation) or a High-Energy Radiation Drain (HERD)Rehammar and Gasparinetti (2023) filter should be implemented.
@@ V SUMMARY AND OUTLOOK
In this review, we have described the requirements for the shielding system to protect the SQC from external IR radiation and stray electromagnetic fields. We have considered a variety of existing shielding systems and compared them to our requirements. We have also considered different techniques for shielding system effectiveness evaluation.
@@ V.1 Recommendations for shielding and filtering
    The shielding system should comprise a minimum of three layers: an IR shield, a superconducting shield, and a ${\mu}$-metal shield. However, it is important to note that simpler shield configurations generally provide better cooling efficiency for the entire system.
@@
    IR shields must be light-tight to prevent the penetration of infrared radiation. The same requirement applies to the sample holder, which serves as the first stage of IR radiation protection. For superconducting and ${\mu}$-metal shields, a closed design is recommended to ensure magnetic field uniformity throughout the shield volume. If this is not feasible, the sample holder should ideally be positioned in the lower third section of the shields.
@@
    To achieve high-efficiency IR shielding, the shield should be constructed with a high-thermal-conductivity base material, such as copper or superconductor-coated copper. For SQC shielding, the sample holder itself may suffice either with an absorbing inner coating (for floating qubits) or without (for other cases). An additional IR shield around the sample holder is advisable if passive measurement circuit elements are nearby. This shield should be positioned as close to the SQC as possible. For the IR-absorbing inner coating of the shield or sample holder, a mixture of Stycast epoxy and $\sim$1 $mm$ SiC particles is recommended.
@@
    Since the sample holder acts as the first shielding layer, its fabrication should follow the same guidelines as those for IR shields (see Recommendations 2-4). The sample holder’s design must be adapted to the SQC configuration and the cryogenic setup’s wiring requirements.
@@
    For optimal magnetic field protection, shielding configurations should be selected based on the required field suppression level for the specific SQC architecture. Regarding single-layer shielding, superconducting shields maintain lower internal magnetic fields compared to $\mu$-metal shields. For double-layer systems, the optimal configuration combines an inner superconducting shield with an outer $\mu$-metal shield, achieving maximum field suppression.
  7. 7.
@@
    To achieve uniform magnetic field distribution within the shielding system, a fully enclosed configuration (including a lid) is recommended. When this is not feasible, the sample holder containing the SQC should be positioned in the lower third section of the shielding volume, where magnetic field is minimal.
@@
    Magnetic components of the cryogenic setup (e.g., isolators, circulators) should be positioned outside the shielding system. Special design features, such as waveguide integration on the shield lid, can further mitigate magnetic field effects on the quantum circuit.
  11. 11.
    If cosmic rays have an undesirable effect on SQC performance, it may be reasonable to place the dilution refrigerator inside a shielded laboratory or in a deep underground location.
  12. 12.
    The best protection of the superconducting quantum circuit from stray electromagnetic fields and IR radiation can be achieved by combining the shielding and microwave filters. Effective filtering of all qubit microwave lines (drive, flux, and readout) requires both microwave and IR filters. For drive lines and readout inputs, either "LPF+IR" or "BPF+IR" filter combinations are recommended, where the LPF or BPF should be selected based on the qubit operating frequency. Eccosorb CR-110 epoxy serves as the optimal IR filter filler for these applications.
    For flux lines, an "LPF+IR" configuration is preferred. The LPF cutoff frequency must be optimized according to the target flux pulse duration, with Eccosorb CR-124 providing superior performance as the absorber for IR filters in this case.
    The readout output line requires similar filtering to the drive and readout input lines ("LPF+IR" or "BPF+IR"). LPF/BPF filters should be matched to the qubit operating frequency, and the IR filter should be minimized in size to reduce output signal attenuation on condition of adequate IR suppression.